# Exploiting Code Symmetries for Learning Program Semantics

### v7

@@ 1 Introduction
Limitations of existing approaches. A popular way to train LLMs to be robust to code symmetries is via large-scale pre-training, where the pre-training dataset likely includes many semantically equivalent code samples (Roziere et al., 2023). While this approach improves the generalization, it is inherently a best-effort attempt and does not guarantee invariance for the pre-trained model. Without the guarantee, the trained model lacks assurance when deployed for program analysis, e.g., the malware can be easily transformed to evade detection (see our evaluation on binary analysis in §6.1). In fact, we find that state-of-the-art LLMs break desired invariances at an alarmingly high rate, e.g., 18% in CodeLlama in predicting function names (Table 1) even for simple code symmetries like a two-statement permutation.
A more direct approach is to explicitly enumerate the code symmetries via data augmentation. However, it is prohibitively expensive due to the sheer number of possible symmetries and their compositions. Similar to pre-training, even exhausting the code symmetries in the augmented training set does not provide any guarantee that the trained model stays invariant to the observed code symmetries. An alternative strategy involves approximating code symmetry structures as priors within the model’s architecture, e.g., control flow graphs, using Graph Neural Nets (GNNs)  (Allamanis et al., 2017). Such approaches are argued to offer a better generalization, e.g., as evidenced in Table 1 where GGNN has the lowest violation rate among the baselines. However, graph architectures often restrict a model’s expressiveness relative to LLM architectures (Ying et al., 2021). Moreover, the existing common practice of encoding code structures into GNNs does not explicitly preserve code symmetries and thus lacks guarantee that the code symmetries are indeed preserved (see §6).
Table 1: Invariance violation rate across different code models (darker colors indicate more violations). Violation SymC (Ours) 0% code2vec 61% code2seq 52% CodeLlama 18% CodeT5 16% DOBF 41% GGNN 7% GPT-4 43% GraphCodeBERT 31% WizardCoder 14%
Refer to caption (a) SymC (Ours) Refer to caption (b) Existing code LLMs
@@
Our approach. In this paper, we investigate how to modify expressive architecture like Transformers to provably impose code semantic priors like code symmetries while still preserving the learning capacity. We introduce a group-theoretic framework to precisely define code symmetries in terms of semantics-preserving statement permutations and create LLM architectures that inherently preserve these symmetries by construction. Using this framework, we present SymC, a variant of LLM architecture designed to provably guarantee invariance to semantics-preserving statement permutations. This is achieved through a $G$-equivariant code representation ($r$) followed by a $G$-invariant prediction ($p$), with $G$ determined based on the graph automorphisms of the code block’s interpretation graph (a generalization of program dependence graph).
@@ 3.1 Invariance & Equivariance for Code Models
Code representation units. We establish formal definitions of the code space as a collection of code blocks, which serves as the input space for representation learning. We then proceed to define representation learning and predictive learning in the code space.
@@
Definition 3.2.
Representation learning for code involves learning a function $r$ that maps a CRU $c\in I^{n}$ to a point in the code representation space $\mathbb{R}^{d\times n}$, $r:I^{n}\rightarrow\mathbb{R}^{d\times n}$ ($d$ denotes the dimension of the embedding to which each instruction is mapped).
Definition 3.3.
Predictive learning for code entails learning a function $p:\mathbb{R}^{d\times n}\rightarrow\mathbb{R}^{L}$ that maps the code representation produced by the representation learning $r$ to a label space $\mathbb{R}^{L}$. $\mathbb{R}^{L}$ becomes concrete in the context of downstream analyses (§5).
In this framework, the earlier layers of the neural network serve as the representation learning function $r$, learning program representations. The subsequent layers serve as the predictive learning function $p$, making predictions based on analysis-specific labels, such as function names. Therefore, the whole network computation can be thought of as a composition of $r$ and $p$, i.e., $p\circ r$.
In the following, we establish formal properties for code analysis models with explicit representation learning $r$ and predictive learning $p$ based on $G$-equivariance/invariance.
@@
Let $G$ be a symmetry group consisting of semantics-preserving transformations applied to a CRU $c\in I^{n}$. A representation function $r:I^{n}\rightarrow\mathbb{R}^{d\times n}$ is $G$-equivariant if for every $g\in G$ and $c\in I^{n}$, we have $g\circ r(c)=r(g\circ c)$.
@@ 3.2 Semantics-Preserving Program Symmetries
Local and global program symmetry. We call $g$ local program symmetry because it acts on a single CRU $c\in I^{n}$. In this paper, we do not consider global program symmetry defined over the entire program space $I^{n}$, e.g., rotation as in image space. This is because each independent sample $c$ will have its own symmetries corresponding to its semantics-preserving transformations, and we develop model architectures that preserve the specific code symmetry group for each individual sample (§4).
@@ 3.4 $Aut(\mathcal{IG})$-Equivariant Code Representation
Biasing self-attention with a distance matrix. To build $Aut(\mathcal{IG})$-equivariant self-attention layers, denoted as $G\!A$, we add a customized distance matrix $d_{\mathcal{IG}}$ to $G\!A$: $G\!A(e)=w_{v}e\cdot(s(w_{k}e^{T}\circ w_{q}e)+d_{\mathcal{IG}})$. Importantly, $d_{\mathcal{IG}}$ should have two properties: (1) $d_{\mathcal{IG}}$ stays invariant when $\sigma\in Aut(\mathcal{IG})$ acts on $\mathcal{IG}$: $d_{\mathcal{IG}}=\sigma(d_{\mathcal{IG}})$, and (2) $d_{\mathcal{IG}}$ commutes with permutation matrix $p_{\sigma}$ ($\sigma\in Aut(\mathcal{IG})$).
We will describe a concrete instantiation of $d_{\mathcal{IG}}$ in §4.2. Based on the two properties, we have the following Theorem (see Appendix B for the proof).
@@
Self-attention $G\!A(e)=w_{v}e\cdot(s(w_{k}e^{T}\cdot w_{q}e)+d_{\mathcal{IG}})$ is $Aut(\mathcal{IG})$-equivariant.
@@ 4.1 Relaxing $\mathcal{IG}$ to Program Dependence Graph
§3.4 demonstrated how to build $Aut(\mathcal{IG})$-equivariant self-attention layers. However, directly constructing $\mathcal{IG}$ is computationally impractical as we need to enumerate all possible execution paths. To address this, we consider program dependence graph (PDG), a sound over-approximation to $\mathcal{IG}$ that explicitly captures the control/data dependencies and can be computed statically and efficiently. PDG ($V_{PDG},E_{PDG}$) is a super graph of $\mathcal{IG}$ due to the conservative data flow analysis, sharing the same vertices but having a superset of edges ($E_{PDG}\supseteq E_{\mathcal{IG}}$). Enforcing PDG to be a super graph of $\mathcal{IG}$ is crucial because a subgraph’s automorphism group is a subgroup of the super graph ($Aut(P\!D\!G)\supseteq Aut(\mathcal{IG})$). Thus, if the self-attention layer is $Aut(P\!D\!G)$-equivariant, it is guaranteed to be $Aut(\mathcal{IG})$-equivariant.
PDG construction. We construct PDG edges based on data/control dependencies between instructions. Three types of data dependencies (read-after-write, write-after-read, and write-after-write) are considered. Control dependencies are included to determine the execution order. These dependencies establish a partial order of instructions, preventing permutations that violate edge directions that might alter the input-output behavior of the program (§3.3).
@@ 4.2 Encoding Graph Structure
We incorporate $d$ into the multi-head self-attention (MHA), ensuring $Aut(P\!D\!G)$-equivariance, and define specific modifications to the attention heads to handle positive and negative distances. Specifically, the first half of the attention heads $M\!H\!A^{i}(e)$, for $i\in[1,h/2]$, are combined with the matrix $dp$ formed by the positive distances in $d$ (denoted as $dp_{ij}=p_{ij}$). The second half of the attention heads $M\!H\!A^{i}(e)$, for $i\in[h/2+1,h]$, are combined with the matrix $dn$ formed by the negative distances in $d$ (denoted as $dn_{ij}=n_{ij}$). The modified attention heads are defined as: (1) $M\!H\!A^{i}(e)=w_{v}e\cdot(s(w_{k}e^{T}\cdot w_{q}e)+dp),i\in[1,h/2]$, (2) $M\!H\!A^{i}(e)=w_{v}e\cdot(s(w_{k}e^{T}\cdot w_{q}e)+dn),i\in[h/2+1,h]$.
@@ 5 Experimental Setup
This section briefly describes the program analysis tasks and code transformations. Due to the space limit, we put the detailed description, e.g., baselines, in Appendix D.
@@ 6.1 Invariance and Generalization
Table 3: Compparing SymC to binary analysis baselines against semantics-preserving permutations. Function signature and memory region prediction are measured in F1. Function similarity detection is measured using AUC (area under the ROC curve). Similar to Table 2, the magnitude of the violation rate is highlighted in red.
@@ 6.2 Training Efficiency
Besides the improved robustness and generalization, SymC is efficient in training in avoiding expensive pre-training efforts, e.g., some may take up to 10 days (Jin et al., 2022). As shown in §6.1, SymC, without any pre-training, outperforms the pre-trained baselines.
This section shows SymC’s performance under the limited training resources. Specifically, we reduce the model sizes and training iterations to test the hypothesis that SymC requires less training effort for similar testing performance due to its improved training efficiency. Figure 6 shows that SymC’s performance (on memory region prediction) remains the highest in both reduced size and training iterations, outperforming PalmTree by 36.9% and 21.4%, respectively. Even in the most strict scenario, SymC remains 38.2% and 15.3% better in both settings.
@@
Figure 6: Comparing SymC and baselines on constrained resources, where we (a) reduce the model weights, and (b) reduce the number of training iterations, and observe how that affects the performance.
@@ 6.3 Ablations
7 Related Work
Code representation learning. Previous research aims to automate software development tasks through code representation learning (Ding et al., 2023; Feng et al., 2020; Guo et al., 2022; Ahmad et al., 2021), employing new architectures and pre-training objectives (Hellendoorn et al., 2019; Allamanis et al., 2017; Sun et al., 2020; Peng et al., 2021; Kim et al., 2021; Guo et al., 2020). However, unlike our approach, these approaches cannot provide any guarantees of encoding semantics.
@@ 7 Related Work
Figure 7: (a) Setting SymC’s self-attention layers invariant in earlier layers. (b) Pre-training SymC with varying pre-training epochs.
@@
Symmetry in machine learning. Symmetry plays a crucial role in creating efficient neural architectures across various domains (Reiser et al., 2022; Wang et al., 2020; Bogatskiy et al., 2020; Perraudin et al., 2019; Cohen & Welling, 2016; Gordon et al., 2019; Dehmamy et al., 2021). Different architectures, such as CNNs, GNNs, and Transformers, leverage symmetry to handle translations, rotations, permutations, etc. (Lee et al., 2019; Cohen & Welling, 2016; Esteves et al., 2018; Hutchinson et al., 2021; Gordon et al., 2019; Romero & Cordonnier, 2020). SymC sets the first step to formalize code semantics learning using symmetry groups.

### v9

@@ 1 Introduction
Limitations of existing approaches. A popular way to train LLMs to be robust to code symmetries is via large-scale pre-training, where the pre-training dataset likely includes many semantically equivalent code samples (Roziere et al., 2023). While this approach improves the generalization, it is inherently a best-effort attempt and does not guarantee invariance for the pre-trained model. Without the guarantee, the trained model lacks assurance when deployed for program analysis (Ullah et al., 2023), e.g., the malware can be easily transformed to evade detection. In fact, we find that state-of-the-art LLMs break desired invariances at an alarmingly high rate, e.g., 18% in CodeLlama in predicting function names (Table 1) even for simple code symmetries like a two-statement permutation.
A more direct approach is to explicitly enumerate the code symmetries via data augmentation. However, it is prohibitively expensive due to the sheer number of possible symmetries and their compositions. Similar to pre-training, even exhausting the code symmetries in the augmented training set does not provide any guarantee that the trained model stays invariant to the observed code symmetries. An alternative strategy involves approximating code symmetry structures as priors within the model’s architecture, e.g., data and control flow graphs, using Graph Neural Nets (GNNs)  (Allamanis et al., 2017). Such approaches are argued to offer a better generalization, e.g., as evidenced in Table 1 where GGNN has the lowest violation rate among the baselines. However, graph architectures often restrict a model’s expressiveness relative to LLM architectures (Ying et al., 2021). Moreover, the existing common practice of encoding code structures into GNNs does not explicitly preserve code symmetries and thus lacks guarantee that the code symmetries are indeed preserved (see §6).
Table 1: Invariance violation rate across different code models (darker colors indicate more violations).
Violation
SymC (Ours) 0%
code2vec 61%
code2seq 52%
CodeLlama 18%
CodeT5 16%
DOBF 41%
GGNN 7%
GPT-4 43%
GraphCodeBERT 31%
WizardCoder 14%
(a) SymC (Ours)
(b) Existing code LLMs
@@
Our approach. In this paper, we investigate how to modify expressive architecture like Transformers to provably impose semantic priors like code symmetries while still preserving the learning capacity. We introduce a group-theoretic framework to precisely define code symmetries in terms of semantics-preserving statement permutations and create LLM architectures that inherently preserve these symmetries by construction. Importantly, such a group-theoretic framework expresses semantic priors in a syntax-agnostic way while being amenable to be encoded in neural architectures (Cohen & Welling, 2016; Romero & Cordonnier, 2020). This makes symmetries an ideal option to represent program semantics as the inductive bias for code LLM architectures.
Using this framework, we present SymC, a variant of LLM architecture designed to provably guarantee invariance to semantics-preserving statement permutations. This is achieved through a $G$-equivariant code representation ($r$) followed by a $G$-invariant prediction ($p$), with $G$ determined based on the graph automorphisms of the code block’s interpretation graph (a generalization of program dependence graph).
@@ 3.1 Invariance & Equivariance for Code Models
Code representation units. We establish formal definitions of the code space as a collection of code blocks, which serve as the input space for representation learning.
@@
We establish formal properties for code analysis models with explicit representation learning $r$ and predictive learning $p$ based on $G$-equivariance/invariance.
@@
Let $G$ be a symmetry group consisting of semantics-preserving transformations applied to a CRU $c\in I^{n}$. A representation function $r:I^{n}\rightarrow\mathbb{R}^{d\times n}$ is $G$-equivariant if for every $g\in G$ and $c\in I^{n}$, we have $g\circ r(c)=r(g\circ c)$ ($d$ denotes the dimension of the embedding to which each instruction is mapped).
@@ 3.4 $Aut(\mathcal{IG})$-Equivariant Code Representation
Biasing self-attention with a distance matrix. To build $Aut(\mathcal{IG})$-equivariant self-attention layers, denoted as $G\!A$, we add a customized distance matrix $d_{\mathcal{IG}}$ to $G\!A$: $G\!A(e)=w_{v}e\cdot s(w_{k}e^{T}\circ w_{q}e+d_{\mathcal{IG}})$. Such a distance matrix is a superset of the adjacency matrix of $\mathcal{IG}$, encoding a richer topology structure of the graph. We relax the definition of distance matrix $d_{\mathcal{IG}}$ here to be no longer symmetrical, as long as it satisfies the following two properties: (1) $d_{\mathcal{IG}}$ stays invariant when $\sigma\in Aut(\mathcal{IG})$ acts on $\mathcal{IG}$: $d_{\mathcal{IG}}=\sigma(d_{\mathcal{IG}})$, and (2) $d_{\mathcal{IG}}$ commutes with permutation matrix $p_{\sigma}$ ($\sigma\in Aut(\mathcal{IG})$).
We will describe a concrete instantiation of $d_{\mathcal{IG}}$ in §4.2. Based on the two properties, we have the following Theorem (with its proof in Appendix B.
@@
Self-attention $G\!A(e)=w_{v}e\cdot s(w_{k}e^{T}\cdot w_{q}e+d_{\mathcal{IG}})$ is $Aut(\mathcal{IG})$-equivariant.
@@ 4.1 Relaxing $\mathcal{IG}$ to Program Dependence Graph
§3.4 demonstrated how to build $Aut(\mathcal{IG})$-equivariant self-attention layers. However, directly constructing $\mathcal{IG}$ is computationally impractical as we need to enumerate all possible execution paths. To address this, we consider program dependence graph (PDG), a sound over-approximation to $\mathcal{IG}$ that explicitly captures the control/data dependencies and can be computed statically and efficiently. PDG ($V_{PDG},E_{PDG}$) is a super graph of $\mathcal{IG}$ due to the conservative data flow analysis, sharing the same vertices but having a superset of edges ($E_{PDG}\supseteq E_{\mathcal{IG}}$). Enforcing PDG to be a super graph of $\mathcal{IG}$ is crucial because a subgraph’s automorphism group is a subgroup of the super graph’s automorphism group ($Aut(P\!D\!G)\supseteq Aut(\mathcal{IG})$). Thus, if the self-attention layer is $Aut(P\!D\!G)$-equivariant, it is guaranteed to be $Aut(\mathcal{IG})$-equivariant.
PDG construction. We construct PDG edges based on data and control dependencies between instructions. We consider three types of data dependencies: read-after-write, write-after-read, and write-after-write. Additionally, control dependencies are included to determine the execution order. These dependencies create a partial order of instructions, preventing permutations that would violate the edge directions and potentially alter the program’s input-output behavior.(§3.3).
To identify the control and data dependencies, we employ a computationally efficient, conservative static analysis approach. For instance, we treat any two statements that access memory as dependent. However, this conservative approach may overlook potential symmetries. We aim to integrate more accurate analyses in the future, such as alias analysis through abstract interpretation or dynamic analysis, to reduce overapproximation while managing the trade-off of increased overhead. Interleaving standard dependency analysis with training/inference to minimize analysis overhead presents a promising direction for future exploration.
Implementing the static analysis to identify symmetries presents an additional challenge, as symmetries are defined at the level of tokens as seen by LLMs, whereas compilers and other program analysis tools typically perform data dependency analysis at the Intermediate Representation (IR) level. It requires significant engineering effort to correlate IR-level dependencies with source-level tokens and subtokens for the self-attention layers. Hence, for our initial prototype, we developed our analysis method based on the source-level ASTs (as described in §D.1) to simplify the mapping from the analysis results back to the source code tokens and subtokens. We note that integrating more sophisticated dependency analysis routines from existing compiler frameworks to identify symmetries is a promising area for future work.
@@ 4.2 Encoding Graph Structure
We incorporate $d$ into the Multi-Head Self-Attention (MHA) to ensure $Aut(P\!D\!G)$-equivariance with specific modifications to the attention heads to handle positive and negative distances. Particularly, the first half of the attention heads $M\!H\!A^{i}(e)$, for $i\in[1,h/2]$, are combined with the matrix $dp$ formed by the positive distances in $d$ (denoted as $dp_{ij}=p_{ij}$). The second half of the attention heads $M\!H\!A^{i}(e)$, for $i\in[h/2+1,h]$, are combined with the matrix $dn$ formed by the negative distances in $d$ (denoted as $dn_{ij}=n_{ij}$). The modified attention heads are defined as: (1) $M\!H\!A^{i}(e)=w_{v}e\cdot s(w_{k}e^{T}\cdot w_{q}e+dp),i\in[1,h/2]$, (2) $M\!H\!A^{i}(e)=w_{v}e\cdot s(w_{k}e^{T}\cdot w_{q}e+dn),i\in[h/2+1,h]$.
@@ 5 Experimental Setup
This section briefly describes the evaluation tasks and code transformations. We put the detailed description, e.g., baselines, evaluation dataset and metrics, etc., in Appendix D.
@@ 6.1 Invariance and Generalization
Table 3: Comparing SymC to binary analysis baselines against semantics-preserving permutations. Function signature and memory region prediction are measured in F1. Function similarity detection is measured using AUC (area under the ROC curve). Similar to Table 2, the magnitude of the violation rate is highlighted in red.
@@ 6.2 Training Efficiency
Besides the improved robustness and generalization, SymC is efficient in avoiding expensive training efforts, e.g., some may take up to 10 days (Jin et al., 2022). As shown in §6.1, SymC, without any pre-training, outperforms the pre-trained baselines.
Therefore, in this section, we aim to study SymC’s performance under the limited training resources. Figure 6 shows that SymC’s performance (on memory region prediction) remains the highest when we reduce the model size and training iterations, outperforming PalmTree by 36.9% and 21.4%, respectively. Even in the most strict scenario, SymC remains 38.2% and 15.3% better in both settings.
@@
Figure 6: Comparing SymC and baselines when we (a) reduce the model weights, and (b) reduce the number of training iterations, and observe how that affects the performance.
@@ 6.3 Ablations
Non-$Aut(P\!D\!G)$-equivariant baselines. We aim to confirm the performance of SymC comes from preserving the symmetry, as opposed to the increased capacity from the distance matrix in the self-attention layers. To test this hypothesis, we compare SymC with the baselines using the same exact architecture, but only changing the distance matrix whose entries are (1) kept exactly the same (i.e., fully permutation equivariant baseline), and (2) derived from the relative distance between the monotonically increasing positions (i.e., non-equivariant relative positional embedding).
@@
Figure 7: (a) Setting SymC’s self-attention layers invariant in earlier layers. (b) Pre-training SymC with varying pre-training epochs. (c) Comparing the $Aut(PDG)$-equivariant layers with the fully permutation equivariant ones (distance matrix with identical values) and non-equivariant ones (default relative position embeddings (Raffel et al., 2020; Wang et al., 2021)).
Figure 7c shows the comparison of testing F1 score by training the three models on memory region prediction tasks. Note that both baselines adopt the same shape of the distance matrix as SymC, so their number of parameters is identical with SymC as well. The results demonstrate that SymC outperforms the fully permutation equivariant one and the non-equivariant one by 14.5% and 36.4%, respectively.
7 Limitations and Future Work
Overhead. SymC incurs additional overhead in generating PDG for each sample, though we have put reasonable engineering effort into making graph construction relatively cheap (Appendix E.2, Figure 11), and we can further optimize the graph construction by interleaving it with the training/inference to further improve the efficiency.
In essence, SymC imposes the equivariance constraints regardless of training. This implies that a slight finetuning to help the model simply adapt to the symmetry constraints while specializing for downstream tasks would be likely enough (Basu et al., 2023). This is exactly why SymC can obtain strong results by training the model from scratch using the typical fine-tuning tasks (not next token prediction or masked language modeling).
Other code symmetries. Our current framework focuses on the permutation group, but it extends to all transformations as long as they form a group. For example, variable name renaming is also a permutation over the entire vocabulary. SymC can also be readily applied to token permutation, e.g., a=a+1 to a=1+a is a symmetry with a proper type inference, e.g., ensuring a is not a string and + is not a string concatenation operator.
However, the expressivity of the code symmetry can be restrictive to certain transformations. For example, insertion and deletion operations (e.g., deadcode elimination) are not invertible. Therefore, a new formalism, i.e., semigroup (Hille & Phillips, 1996) which relaxes the requirement of invertibility (§2), is required. Moreover, it is often an expensive manual effort to identify the symmetry group structure for arbitrary code transformations. For example, it is unclear how a compiler optimization pass, e.g., gcc O3, can be expressed using a sequence of insertion, deletion, and permutation operations. It would be an interesting future direction to uncover the unknown code symmetry from these transformations (Huh, 2024).
8 Related Work
Code representation learning. Previous research aims to automate software development tasks through code representation learning (Ding et al., 2023; Feng et al., 2020; Guo et al., 2022; Ahmad et al., 2021). Typical methodologies involve employing new model architectures and pre-training objectives with the goal of learning code representation that can be reused for other downstream program analysis and understanding tasks (Hellendoorn et al., 2019; Bieber et al., 2020; Pei et al., 2022, 2021; Allamanis et al., 2017; Sun et al., 2020; Peng et al., 2021; Kim et al., 2021; Guo et al., 2020). However, unlike our approach, these approaches do not provide any guarantees that the underlying models can be robust against any semantics-preserving transformations including simple permutations.
An alternative popular setting is to ground the program with its execution behavior so it learns semantics-aware code representations (Wen et al., 2024; Pei et al., 2022; Ding et al., 2023; Nye et al., 2021; Souza & Pradel, 2023; Wang & Su, 2020; Bieber et al., 2022; Ye et al., 2022). One caveat of incorporating execution in SymC is that the dynamic traces are often incomplete to expose all possible code behaviors. Therefore, it might suffer from identifying false symmetries. However, as conservative static analysis can miss many true symmetries, dynamic traces can assist the analysis in canonicalizing transformations that are hard to reason about statically. In our current formulation based on PDG, we only consider symmetries for all input, but it can be relaxed to a subset of inputs, which is an exciting future work.
@@ 8 Related Work
Symmetry in deep learning. Symmetry plays a crucial role in creating efficient neural architectures in various domains (Reiser et al., 2022; Wang et al., 2020; Bogatskiy et al., 2020; Perraudin et al., 2019; Cohen & Welling, 2016; Gordon et al., 2019; Dehmamy et al., 2021). Different architectures, such as CNNs, GNNs, and Transformers, leverage symmetry as the inductive bias to build model architectures robust to geometric transformations such as translations, rotations, permutations, etc. (Lee et al., 2019; Cohen & Welling, 2016; Esteves et al., 2018; Hutchinson et al., 2021; Gordon et al., 2019; Romero & Cordonnier, 2020). SymC sets the first step to formalize code semantics learning using symmetry groups.
Recent studies have explored various strategies beyond symmetry groups to encode the geometric properties of input, aiming to enhance the model’s robustness, generalization, and sample efficiency. These efforts include integrating symbolic operations into the neural architecture (Chaudhuri et al., 2021; Li et al., 2023; Kim et al., 2017; Hu et al., 2016) and designing specialized loss functions (Fort et al., 2019; Basu et al., 2023; Huh, 2024). Neural-symbolic approaches are particularly appealing for modeling programs, as programs typically encompass symbolic elements, such as syntactic and semantic rules, and fuzzy elements, such as natural language comments and function/variable names. However, developing differentiable components for symbolic rules might require extensive effort from the domain expert (Yi et al., 2018; Garcez et al., 2022). SymC represents a new alternative to express program semantics in a way that is amenable to engineering the neural architecture, thanks to the rich literature in the domain of geometric deep learning (Bronstein et al., 2017). By representing program semantics as code symmetry groups, our approach holds the promise of enhancing inference efficiency by bypassing the interpretation of symbolic rules, which often face scalability challenges in traditional program analysis.